find A x X for the following vectors :

За, + 4а, В (b) A a, + 2a, + 4a, B (c) A = 2a, +5a, + la, B 5a, 1a, (a) A = 2a, 2a,+6a, a, + 3a,

electromagnetism discipline.
Answer #1

& et A = a, i thitik and then. B = azit bai tak AXB = 1 | qz bz (2) A = 29x-zayttar, B=5ay- 19. x ay az AXB = 41 is a lo 5 -b) A = Apr 24 +447. B292460 199 947 AB = 24- 206 : 7/12-6) - 90 (6-9) +91(0-4) - 129p + 2 - 491 AxB = 12 + 2 4-497 (c) A = 29= A B 5 - 1547 و 5 - : ;۸b -وو - و دا، (او -) - ه ( و ) + أم (ه -5)